# ZMP-HH-25/5

### v1

@@ 6.2 Multi-Component Type II Degenerations
where the end-components $S_{1}$ and $S_{n}$ are rational surfaces. The components $S_{a}$ intersect in such a way that only $C_{a,a+1}$ are non-vanishing. The $C_{a,a+1}$ are elliptic curves sharing the same complex structure. For K3 surfaces this condition on type II degenerations is already manifest at the level of the mixed Hodge structure. To see this, consider the Hodge–Deligne diamond for a K3 surface, which is defined in complete analogy to the Hodge–Deligne diamond for threefolds. For type II degenerations of K3 surfaces the Hodge–Deligne diamond is depicted in Figure -953. Here the spaces represented by red dots are one-dimensional. The graded space $\text{Gr}_{1}$ can hence be identified with the pure Hodge structure on the middle cohomology of a torus. In analogy to the Tyurin degenerations, this torus can be identified with the curves $C_{a,a+1}$ arising at the intersections of the irreducible components of $S_{0}$. The space $\text{Gr}_{1}$ being isomorphic to the Hodge structure of a single torus further explains why all curves $C_{a,a+1}$ have the same complex structure. In the case of K3 surfaces, the limiting mixed Hodge structure associated with a type II limit fixes the possible kind of degenerate geometries of the form (6.4) that can arise in these limits.
@@
Figure -953: Hodge-Deligne diamond for a type II degeneration of a K3 surface, with the populated horizontal rows representing the spaces ${\rm Gr}_{1}$, ${\rm Gr}_{2}$, ${\rm Gr}_{3}$.
@@
We now return to general type II limits for Calabi–Yau threefolds. Based on our discussion at the end of Section 3.2, from the perspective of low-energy theory general type II limits are expected to correspond to emergent string limits. The Hodge–Deligne diamond for such limits is depicted in Figure -958. As reviewed in Section 2.1, the Calabi–Yau condition for the underlying threefold $V$ constrains $i^{2,0}=i^{0,2}=1$. Therefore, in a generic type II limit we obtain the mixed Hodge structure for a Kähler surface with a unique holomorphic $(2,0)$-form, which can either be a K3 or an Abelian surface. If we consider a semi-stable degeneration of the kind in (6.1) this implies that at least one of the $M_{i,j}$ is a K3 or Abelian surface with a pure Hodge structure contained in $\text{Gr}_{2}$. If there are multiple $M_{i,j}$ of this kind, they all have to be describable by the same Hodge structure — otherwise $i^{2,0}$ would be larger than $1$ because each surface would contribute an independent $(2,0)$ form. Therefore these $M_{i,j}$ are either all Abelian surfaces with the same complex structure or all K3 surfaces that are polarised by the same lattice $\Lambda_{\rm pol}$. However, unlike for type II Kulikov models, from the algebraic perspective this does not yet exclude the existence of further non-zero $M_{i,j}$ corresponding to Fano surfaces with $h^{2,0}=h^{0,2}=0$. The reason is that the Hodge theory does not constrain the dimension of $I^{1,1}$, i.e. the dimension of the central red dot in Figure -958, which is the free parameter $b={i}^{1,1}$ characterising the IIb degeneration.
@@
Since at least one of the $M_{i_{0},j_{0}}$ has to be a K3 surface, we get $\text{dim}\,H^{2}(M_{i_{0},j_{0}})=22$ scalar fields, out of which three are right-moving and 19 are left-moving, or in the case of $M_{i_{0},j_{0}}$ being an Abelian surface $\text{dim}\,H^{2}(M_{a_{0},b_{0}})=6$ scalar fields (three left- and three right-moving) and $82b_{1}(M_{i_{0},j_{0}})=$ 8 left-moving fermions. If there was another, inequivalent surface $M_{i_{1},j_{1}}$ additional localised zero modes from $H^{2}(M_{i_{1},j_{1}})$ would contribute to the worldsheet spectrum of the string. Furthermore, also upon reduction of $B_{6}$ and $C_{6}$ along $H^{4}(M_{i_{1},j_{1}})\neq 0$ we would obtain additional massless worldsheet degrees of freedom. Hence, also taking into account the additional zero modes discussed in Section 4 gives a string worldsheet with right-moving central charge
@@
for which the non-vanishing double surfaces $M_{i,j}=X_{i}\cap X_{j}$ are either all Abelian with the same complex structure or all K3 surfaces that are polarised by the same lattice $\Lambda_{\rm pol}$ of rank $(1,19-b)$. In particular, the parameter $b$ characterzing the IIb degeneration of the family is bounded as
@@ 7 Discussion and Conclusions
[…]
@@ Acknowledgements
We thank Charles Doran, Thomas Grimm, Amir Kashani-Poor, Albrecht Klemm, Seung-Joo Lee, Boris Pioline, and Jörg Teschner for discussions. The work of B.F. is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). The work of J.M., T.W., and M.W. is supported in part by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2121 Quantum Universe 390833306, by Deutsche Forschungsgemeinschaft through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland” and by Deutsche Forschungsgemeinschaft through the Collaborative Research Center 1624 “Higher Structures, Moduli Spaces and Integrability.”
@@ Appendix A Limiting mixed Hodge Structures
such that the following two conditions are satisfied [^42] Although we will not make use of it directly, let us record the explicit formula of the weight filtration in terms of images and kernels of $N$ $W_{\ell+3}(N)=\sum_{j\geq\mathrm{max}{(-1,\ell)}}\mathrm{ker}\,N^{j+1}\cap\mathrm{im}\,N^{j-\ell}\,.$ (A.4)
@@
such that the various $(p,q)$-components satisfy a certain relation under complex conjugation. [^43] The precise relation is $\overline{I^{p,q}}=I^{q,p}\quad\mathrm{mod}\quad\bigoplus_{r<q,s<p}I^{r,s}\,.$ (A.8) Again, let us emphasise that this decomposition can be different for each singularity $\Delta_{k_{1}\cdots k_{r}}$, such that one obtains in total $12^{h^{2,1}}-$ 1 different limiting mixed Hodge structures, within the local patch $u^{1},\ldots,u^{h^{2,1}}=0$ in $\mathcal{M}_{\mathrm{V}}$ we are currently considering. The relation between the weight filtration $W_{\ell}$ and the Deligne splitting $I^{p,q}$ is given by
@@
$\mathrm{I}_{0}\stackrel{{\scriptstyle u^{k_{1}}\to 0}}{{\to}}I^{p,q}(\Delta_{k_{1}})\to\cdots\stackrel{{\scriptstyle u^{k_{n}}\to 0}}{{\to}}I^{p,q}(\Delta_{k_{1}\cdots k_{n}})\,,$ (A.15)
@@
$\mathrm{I}_{0}\stackrel{{\scriptstyle u_{1}\to 0}}{{\to}}I^{p,q}(\Delta_{1})\stackrel{{\scriptstyle u_{2}\to 0}}{{\to}}I^{p,q}(\Delta_{12})\,,$ (A.16)
$\mathrm{I}_{0}\stackrel{{\scriptstyle u_{2}\to 0}}{{\to}}I^{p,q}(\Delta_{2})\stackrel{{\scriptstyle u_{1}\to 0}}{{\to}}I^{p,q}(\Delta_{21})\,.$ (A.17)
@@ Appendix B The Mayer–Vietoris exact Sequence
$\cdots\stackrel{{\scriptstyle\delta_{k-1}}}{{\to}}H^{k}(V_{0})\stackrel{{\scriptstyle\alpha_{k}}}{{\to}}H^{k}(X_{1})\oplus H^{k}(X_{2})\stackrel{{\scriptstyle\beta_{k}}}{{\to}}H^{k}(Z)\stackrel{{\scriptstyle\delta_{k}}}{{\to}}H^{k+1}(V_{0})\to\cdots\,.$ (B.2)
@@
    In particular, note that $\eta$ is closed, so that it gives rise to a cohomology class $[\eta]$, but it is not (globally) exact. Alternatively, one may write
@@ C.1 Mixed Hodge Structures from Mayer–Vietoris
Consider the case where, instead of a Tyurin degeneration, $V_{0}$ is a genus three curve that acquires two nodal singularities, as illustrated in Figure -952.
Figure -952: A degeneration of a genus three curve that acquires two nodal singularities. The 1-cycles indicated in red are examples of 1-cycles in $H_{1}(X_{1})\oplus H_{1}(X_{2})$, while the 1-cycle indicated in blue arises from the combination of two 1-chains on $X_{1}$ and $X_{2}$ combining at the intersection points to give rise to a 1-cycle on the total space $V_{0}$.
@@
where each $X_{i}$ is a genus one curve, and the intersection $X_{1}\cap X_{2}=Z$ is given by two points. Our goal is to characterise $H^{1}(V_{0})$. To do so, let us first discuss the homology $H_{1}(V_{0})$ instead. The picture in Figure -952 suggests that there are two kinds of 1-cycles. First, there are the 1-cycles coming from $H_{1}(X_{i})$ separately, some of which are indicated in red. Second, there is the 1-cycle indicated in blue which consists of two 1-chains on $X_{1}$ and $X_{2}$ that meet at the intersection points. This intuitive picture is made precise by the Mayer–Vietoris exact sequence for homology, the relevant part of which is given by
$0\to H_{1}(X_{1})\oplus H_{1}(X_{2})\to H_{1}(V_{0})\stackrel{{\scriptstyle\partial_{*}}}{{\to}}H_{0}(Z)\to\dots\,.$ (C.2)
@@
$0\to H^{0}(Z)\stackrel{{\scriptstyle d^{*}}}{{\to}}H^{1}(V_{0})\stackrel{{\scriptstyle(k^{*},l^{*})}}{{\to}}H^{1}(X_{1})\oplus H^{1}(X_{2})\to\cdots\,.$ (C.3)
@@
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}H^{1}(V_{0})/\mathrm{ker}(k^{*},l^{*})\,,$
$\stackrel{{\scriptstyle\mathrm{(b)}}}{{\cong}}\mathrm{im}(k^{*},l^{*})\,,$
@@
$\stackrel{{\scriptstyle\mathrm{(a)}}}{{\cong}}H^{0}(Z)/\mathrm{ker}\,d^{*}\,,$
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}H^{0}(Z)\,,$
@@
$\cdots\to H^{2}(X_{1})\oplus H^{2}(X_{2})\stackrel{{\scriptstyle i^{*}-j^{*}}}{{\to}}H^{2}(Z)\stackrel{{\scriptstyle d^{*}}}{{\to}}H^{3}(V_{0})\stackrel{{\scriptstyle(k^{*},l^{*})}}{{\to}}H^{3}(X_{1})\oplus H^{3}(X_{2})\to 0\,.$ (C.7)
@@
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}H^{3}(V_{0})/\mathrm{ker}(k^{*},l^{*})\,,$
$\stackrel{{\scriptstyle\mathrm{(b)}}}{{\cong}}\mathrm{im}(k^{*},l^{*})\,,$
=$\stackrel{{\scriptstyle\mathrm{(c)}}}{{=}}H^{3}(X_{1})\oplus H^{3}(X_{2})\,,$
@@
$\stackrel{{\scriptstyle\mathrm{(a)}}}{{\cong}}H^{2}(Z)/\mathrm{ker}\,d^{*}\,,$
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}H^{2}(Z)/\mathrm{im}(i^{*}-j^{*})\,,$
=$\stackrel{{\scriptstyle\mathrm{(c)}}}{{=}}H^{2}(Z)/(\mathrm{im}(i^{*})+\mathrm{im}(j^{*}))\,,$

### v2

@@ 6.2 Multi-Component Type II Degenerations
where the end-components $S_{1}$ and $S_{n}$ are rational surfaces. The components $S_{a}$ intersect in such a way that only $C_{a,a+1}$ are non-vanishing. The $C_{a,a+1}$ are elliptic curves sharing the same complex structure. For K3 surfaces this condition on type II degenerations is already manifest at the level of the mixed Hodge structure. To see this, consider the Hodge–Deligne diamond for a K3 surface, which is defined in complete analogy to the Hodge–Deligne diamond for threefolds. For type II degenerations of K3 surfaces the Hodge–Deligne diamond is depicted in Figure 7. Here the spaces represented by red dots are one-dimensional. The graded space $\text{Gr}_{1}$ can hence be identified with the pure Hodge structure on the middle cohomology of a torus. In analogy to the Tyurin degenerations, this torus can be identified with the curves $C_{a,a+1}$ arising at the intersections of the irreducible components of $S_{0}$. The space $\text{Gr}_{1}$ being isomorphic to the Hodge structure of a single torus further explains why all curves $C_{a,a+1}$ have the same complex structure. In the case of K3 surfaces, the limiting mixed Hodge structure associated with a type II limit fixes the possible kind of degenerate geometries of the form (6.4) that can arise in these limits.
@@
Figure 7: Hodge-Deligne diamond for a type II degeneration of a K3 surface, with the populated horizontal rows representing the spaces ${\rm Gr}_{1}$, ${\rm Gr}_{2}$, ${\rm Gr}_{3}$.
@@
We now return to general type II limits for Calabi–Yau threefolds. Based on our discussion at the end of Section 3.2, from the perspective of low-energy theory general type II limits are expected to correspond to emergent string limits. The Hodge–Deligne diamond for such limits is depicted in Figure 2. As reviewed in Section 2.1, the Calabi–Yau condition for the underlying threefold $V$ constrains $i^{2,0}=i^{0,2}=1$. Therefore, in a generic type II limit we obtain the mixed Hodge structure for a Kähler surface with a unique holomorphic $(2,0)$-form, which can either be a K3 or an Abelian surface. If we consider a semi-stable degeneration of the kind in (6.1) this implies that at least one of the $M_{i,j}$ is a K3 or Abelian surface with a pure Hodge structure contained in $\text{Gr}_{2}$. If there are multiple $M_{i,j}$ of this kind, they all have to be describable by the same Hodge structure — otherwise $i^{2,0}$ would be larger than $1$ because each surface would contribute an independent $(2,0)$ form. Therefore these $M_{i,j}$ are either all Abelian surfaces with the same complex structure or all K3 surfaces that are polarised by the same lattice $\Lambda_{\rm pol}$. However, unlike for type II Kulikov models, from the algebraic perspective this does not yet exclude the existence of further non-zero $M_{i,j}$ corresponding to Fano surfaces with $h^{2,0}=h^{0,2}=0$. The reason is that the Hodge theory does not constrain the dimension of $I^{1,1}$, i.e. the dimension of the central red dot in Figure 2, which is the free parameter $b={i}^{1,1}$ characterising the IIb degeneration.
@@
Since at least one of the $M_{i_{0},j_{0}}$ has to be a K3 surface, we get $\text{dim}\,H^{2}(M_{i_{0},j_{0}})=22$ scalar fields, out of which three are right-moving and 19 are left-moving, or in the case of $M_{i_{0},j_{0}}$ being an Abelian surface $\text{dim}\,H^{2}(M_{a_{0},b_{0}})=6$ scalar fields (three left- and three right-moving) and $2b_{1}(M_{i_{0},j_{0}})=8$ left-moving fermions. If there was another, inequivalent surface $M_{i_{1},j_{1}}$ additional localised zero modes from $H^{2}(M_{i_{1},j_{1}})$ would contribute to the worldsheet spectrum of the string. Furthermore, also upon reduction of $B_{6}$ and $C_{6}$ along $H^{4}(M_{i_{1},j_{1}})\neq 0$ we would obtain additional massless worldsheet degrees of freedom. Hence, also taking into account the additional zero modes discussed in Section 4 gives a string worldsheet with right-moving central charge
@@
for which the non-vanishing double surfaces $M_{i,j}=X_{i}\cap X_{j}$ are either all Abelian with the same complex structure or all K3 surfaces that are polarised by the same lattice $\Lambda_{\rm pol}$ of rank (1,19b)(1,19-b). In particular, the parameter $b$ characterzing the IIb degeneration of the family is bounded as
@@ 7 Discussion and Conclusions
Even though the criticality of the string ensures the existence of an infinite tower of states in its tensionless limit, its excitation spectrum depends on the details of the worldsheet theory. This is particularly relevant in the context of the Weak Gravity Conjecture [98], for which string excitations play a significant role in weak coupling limits [9, 88, 99, 100, 23, 101, 102, 103, 92, 104]. To determine the spectrum of charged string excitations, we have to know the current algebras realised on the string worldsheet. In Section 4 we have determined the global worldsheet algebras by studying the interactions on the string. As we have seen, the free modes are associated with elements in the transcendental lattice, $\Lambda_{\rm trans}$, of the K3 arising at the Tyurin degenerations. The resulting global algebra on the string worldsheet gives rise precisely to the gauge theory in the bulk that becomes weakly coupled in the emergent string limit, as expected. Based on the study of the kinetic terms for the scalar modes on the string, for $4\leq\text{rk}\,\Lambda_{\rm trans}\leq 20$ we have identified the string as the heterotic string on ${\rm K3}_{\rm het}\times T^{2}$, with a perturbative heterotic gauge algebra of rank $\text{rk}\,\Lambda_{\rm trans}$. In the other cases, we have indicated a possible realisation of the heterotic duality, but due to left-/right-moving asymmetry, the exact details of these models remain to be understood further.
@@ Acknowledgements
We thank Charles Doran, Thomas Grimm, Amir Kashani-Poor, Albrecht Klemm, Seung-Joo Lee, Boris Pioline, and Jörg Teschner for discussions. The work of B.H. is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). The work of J.M., T.W., and M.W. is supported in part by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2121 Quantum Universe 390833306, by Deutsche Forschungsgemeinschaft through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland” and by Deutsche Forschungsgemeinschaft through the Collaborative Research Center 1624 “Higher Structures, Moduli Spaces and Integrability.”
@@ Appendix A Limiting mixed Hodge Structures
such that the following two conditions are satisfied [^43] Although we will not make use of it directly, let us record the explicit formula of the weight filtration in terms of images and kernels of NN $W_{\ell+3}(N)=\sum_{j\geq\mathrm{max}{(-1,\ell)}}\mathrm{ker}\,N^{j+1}\cap\mathrm{im}\,N^{j-\ell}\,.$ (A.4)
@@
such that the various $(p,q)$-components satisfy a certain relation under complex conjugation. [^44] The precise relation is $\overline{I^{p,q}}=I^{q,p}\quad\mathrm{mod}\quad\bigoplus_{r<q,s<p}I^{r,s}\,.$ (A.8) Again, let us emphasise that this decomposition can be different for each singularity $\Delta_{k_{1}\cdots k_{r}}$, such that one obtains in total $2^{h^{2,1}}-1$ different limiting mixed Hodge structures, within the local patch $u^{1},\ldots,u^{h^{2,1}}=0$ in $\mathcal{M}_{\mathrm{V}}$ we are currently considering. The relation between the weight filtration $W_{\ell}$ and the Deligne splitting $I^{p,q}$ is given by
@@
$\mathrm{I}_{0}\mathrel{{\mathop{\to}\limits^{u^{k_{1}}\to 0}}}I^{p,q}(\Delta_{k_{1}})\to\cdots\mathrel{{\mathop{\to}\limits^{u^{k_{n}}\to 0}}}I^{p,q}(\Delta_{k_{1}\cdots k_{n}})\,$, (A.15)
@@
$\mathrm{I}_{0}\mathrel{{\mathop{\to}\limits^{u_{1}\to 0}}}I^{p,q}(\Delta_{1})\mathrel{{\mathop{\to}\limits^{u_{2}\to 0}}}I^{p,q}(\Delta_{12})\,$, (A.16)
$\mathrm{I}_{0}\mathrel{{\mathop{\to}\limits^{u_{2}\to 0}}}I^{p,q}(\Delta_{2})\mathrel{{\mathop{\to}\limits^{u_{1}\to 0}}}I^{p,q}(\Delta_{21})\,.$ (A.17)
@@ Appendix B The Mayer–Vietoris exact Sequence
$\cdots\mathrel{{\mathop{\to}\limits^{\delta_{k-1}}}}H^{k}(V_{0})\mathrel{{\mathop{\to}\limits^{\alpha_{k}}}}H^{k}(X_{1})\oplus H^{k}(X_{2})\mathrel{{\mathop{\to}\limits^{\beta_{k}}}}H^{k}(Z)\mathrel{{\mathop{\to}\limits^{\delta_{k}}}}H^{k+1}(V_{0})\to\cdots\,.$ (B.2)
@@
    In particular, note that $\eta$ is closed, so that it gives rise to a cohomology class $\eta]$, but it is not (globally) exact. Alternatively, one may write
@@ C.1 Mixed Hodge Structures from Mayer–Vietoris
Consider the case where, instead of a Tyurin degeneration, $V_{0}$ is a genus three curve that acquires two nodal singularities, as illustrated in Figure 8.
Figure 8: A degeneration of a genus three curve that acquires two nodal singularities. The 1-cycles indicated in red are examples of 1-cycles in $H_{1}(X_{1})\oplus H_{1}(X_{2})$, while the 1-cycle indicated in blue arises from the combination of two 1-chains on $X_{1}$ and $X_{2}$ combining at the intersection points to give rise to a 1-cycle on the total space $V_{0}$.
@@
where each $X_{i}$ is a genus one curve, and the intersection $X_{1}\cap X_{2}=Z$ is given by two points. Our goal is to characterise $H^{1}(V_{0})$. To do so, let us first discuss the homology $H_{1}(V_{0})$ instead. The picture in Figure 8 suggests that there are two kinds of 1-cycles. First, there are the 1-cycles coming from $H_{1}(X_{i})$ separately, some of which are indicated in red. Second, there is the 1-cycle indicated in blue which consists of two 1-chains on $X_{1}$ and $X_{2}$ that meet at the intersection points. This intuitive picture is made precise by the Mayer–Vietoris exact sequence for homology, the relevant part of which is given by
$0\to H_{1}(X_{1})\oplus H_{1}(X_{2})\to H_{1}(V_{0})\mathrel{{\mathop{\to}\limits^{\partial_{*}}}}H_{0}(Z)\to\dots\,.$ (C.2)
@@
$0\to H^{0}(Z)\mathrel{{\mathop{\to}\limits^{d^{*}}}}H^{1}(V_{0})\mathrel{{\mathop{\to}\limits^{(k^{*},l^{*})}}}H^{1}(X_{1})\oplus H^{1}(X_{2})\to\cdots\,.$ (C.3)
@@
$\mathrel{{\mathop{=}\limits^{\mathrm{(a)}}}}H^{1}(V_{0})/\mathrm{ker}(k^{*},l^{*})\,$,
$\mathrel{{\mathop{\cong}\limits^{\mathrm{(b)}}}}\mathrm{im}(k^{*},l^{*})\,$,
@@
$\mathrel{{\mathop{\cong}\limits^{\mathrm{(a)}}}}H^{0}(Z)/\mathrm{ker}\,d^{*}\,$,
$\mathrel{{\mathop{=}\limits^{\mathrm{(b)}}}}H^{0}(Z)\,$,
@@
$\cdots\to H^{2}(X_{1})\oplus H^{2}(X_{2})\mathrel{{\mathop{\to}\limits^{i^{*}-j^{*}}}}H^{2}(Z)\mathrel{{\mathop{\to}\limits^{d^{*}}}}H^{3}(V_{0})\mathrel{{\mathop{\to}\limits^{(k^{*},l^{*})}}}H^{3}(X_{1})\oplus H^{3}(X_{2})\to 0\,.$ (C.7)
@@
$\mathrel{{\mathop{=}\limits^{\mathrm{(a)}}}}H^{3}(V_{0})/\mathrm{ker}(k^{*},l^{*})\,$,
$\mathrel{{\mathop{\cong}\limits^{\mathrm{(b)}}}}\mathrm{im}(k^{*},l^{*})\,$,
$\mathrel{{\mathop{=}\limits^{\mathrm{(c)}}}}H^{3}(X_{1})\oplus H^{3}(X_{2})\,$,
@@
$\mathrel{{\mathop{\cong}\limits^{\mathrm{(a)}}}}H^{2}(Z)/\mathrm{ker}\,d^{*}\,$,
$\mathrel{{\mathop{=}\limits^{\mathrm{(b)}}}}H^{2}(Z)/\mathrm{im}(i^{*}-j^{*})\,$,
$\mathrel{{\mathop{=}\limits^{\mathrm{(c)}}}}H^{2}(Z)/(\mathrm{im}(i^{*})+\mathrm{im}(j^{*}))\,$,